OpenStudy (anonymous):

Condense the logarithm 12. Log 10 - Log 5 13. Log x - Log 9

zepdrix (zepdrix):

\[\large\rm \log(a)-\log(b)=\log\left(\frac{a}{b}\right)\]

OpenStudy (anonymous):

so is the first one log10/ log 5= .301?

zepdrix (zepdrix):

\(\rm log(2)\approx 0.301\) yes :)

OpenStudy (anonymous):

Thanks! How do i do the second one?

OpenStudy (anonymous):

Since there is an x

zepdrix (zepdrix):

For the first one, notice that we could have stopped here,\[\large\rm \log(10)-\log(5)\quad=\log\left(\frac{10}{5}\right)\quad=\log(2)\]So for the other question, that's where you'll stop after applying the rule. No calculator step.

OpenStudy (anonymous):

so log x/ log 9 = log (9)? idk

zepdrix (zepdrix):

Oh sorry :) I guess we're stopping even one step earlier than I realized,\[\large\rm \log(x)-\log(9)\quad=\log\left(\frac{x}{9}\right) \qquad\qquad \color{green}{\checkmark}\]Notice that when we use this rule, `we don't divide logs`. We instead `divide the insides`. You should not end up with two logs! :)

OpenStudy (anonymous):

ohhh okay (: can you help me with two more questions!! your my only hope because i have to leave soon and none else can help me with this